Find three smallest consecutive natural numbers such that the difference between one - third of the largest and one - fifth of the smallest is atleast 3.
Answer by htmentor(1343)

Let the numbers be n, n+1 and n+2
(1/3)(n+2) - (1/5)n >= 3
n/3 + 2/3 - n/5 >= 3
2n/15 >= 7/3
2n >= 35
n >= 17 1/2
The next integer greater than 17 1/2 is 18
Thus the three smallest numbers satisfying the condition are 18, 19 and 20